Solve the following systems by substitution. Example 2 (p 712) and Example 4 (p 713) are the key models to go by for solving these exercises….
16) 5x – 2y = -5
y - 5x = 3

Answer by checkley75(3666)

5x-2y=-5
-5x+y=3
-----------------------
-y=-2
y=2 answer.
5x-2*2=-5
5x-4=-5
5x=-5+4
5x=-1
x=-1/5 answer.
proof
-5*-1/5+2=3
5/5+2=3
1+2=3
3=3